Use resonance structures to identify the areas of high and low electron density in the following compounds: (a) CH3COH (b) CH3CONH2 (c) CH3CNHH (d) CH3COOCH3 (e) CH3OCONH2 (f)NHO (g)CNNH2 (h) OO (i)ONH (j)HCOCH3O
Answer #1

(a) - H CH OS (b) C CHCNH C -NH NH (c) CH -H C -CH CHC-CH3 = CH-c-c Cd CH :dp CH-O-C-NH, (e) CH HN=O HN-O NEC-NH N=C = NH NEC